OpenStudy (error1603):

Solve 7(x + 8) + 2(x – 4) = 12?

OpenStudy (error1603):

sooo

OpenStudy (error1603):

No one

OpenStudy (er.mohd.amir):

7x+56+2x-8=12 9x=12-56+8 9x=-36 x=-4

OpenStudy (error1603):

I was getting X=7 what am i doing wrong?